Simultaneous equation
Exercise; substitution method 1
Use the substitution method to solve the following:
Challenge:
5x+3y=15
2x-3y=-1 <= hint make x the subject
I have stuck. This is the last question I can't solve? Hicc

Answer by Cromlix(4381)

5x + 3y = 15
2x - 3y = -1
Add
7x = 14
x = 2
Substitute x = 2
in 2x - 3y = -1
2(2) - 3y = -1
4 - 3y = -1
- 3y = -1 - 4
- 3y = - 5
y = -5/-3
y = 5/3
x = 2 y = 5/3